OpenStudy (anonymous):

How many solutions does the equation have? A+5=1/5 (15+5a) A.one solution B.infinite solutions C.no solution

OpenStudy (chosenmatt):

B again

OpenStudy (confluxepic):

The equation has one solution.

OpenStudy (shinalcantara):

a + 5 = 1/5(15 + 5a) a + 5 = 3 + a a-a =5-3 0 is not equal to 2 then there's no solution

OpenStudy (anonymous):

thank you @shinalcantara

OpenStudy (shinalcantara):

yw @amber_ remember that when it would have the same situation as that, there's no solution and those that are having equal values left and right of the equation are having infinite solutions. say for example: 2(2x + 1) = 4x + 2 4x + 2 = 4x + 2 -----notice that the left and right sides are equal making the equation, 0=0, then this type has infinite solutions that whatever values of x you will plug in will be true to the equation